# Salient Region-Guided Spacecraft Image Arbitrary-Scale Super-Resolution Network

### v1

@@ Abstract
Spacecraft image super-resolution seeks to enhance low-resolution spacecraft images into high-resolution ones. Although existing arbitrary-scale super-resolution methods perform well on general images, they tend to overlook the difference in features between the spacecraft core region and the large black space background, introducing irrelevant noise. In this paper, we propose a salient region-guided spacecraft image arbitrary-scale super-resolution network (SGSASR), which uses features from the spacecraft core salient regions to guide latent modulation and achieve arbitrary-scale super-resolution. Specifically, we design a spacecraft core region recognition block (SCRRB) that identifies the core salient regions in spacecraft images using a pre-trained saliency detection model. Furthermore, we present an adaptive-weighted feature fusion enhancement mechanism (AFFEM) to selectively aggregate the spacecraft core region features with general image features by dynamic weight parameter to enhance the response of the core salient regions. Experimental results demonstrate that the proposed SGSASR outperforms state-of-the-art approaches.
Index Terms:
@@
Spacecraft image, Arbitrary-scale super-resolution, Latent modulation, Salient region-guided
I Introduction
@@
Spacecraft image is the space target image obtained through radar or optical technology. The clear and high-quality spacecraft image is the fundamental assurance for successfully completing many space missions such as spacecraft pose estimation [1, 2, 3], spacecraft control [4, 5, 6] and spacecraft 3D reconstruction [7, 8, 9]. However, due to the large imaging distance, the harsh space lighting environment and sensor interference, it is difficult to obtain high-resolution (HR) spacecraft image [10].
@@
Spacecraft image super-resolution (SR) aims to restore HR spacecraft images from low-resolution (LR) spacecraft images. Although many spacecraft image SR methods have been proposed [11, 12], they typically learn upsampling functions for a specific scale factor and require training different models for different SR scales [13]. When there is a lack of suitable scale training data or the degradation scale of images is unfixed, these methods yield unsatisfactory performance. In practical aerospace application scenarios, it is relatively difficult to obtain real spacecraft HR images to construct training datasets. The degradation scale of different images in the same spacecraft image dataset is also unfixed because of the complex space environmental conditions. In addition, storing a large number of models with different super-resolution scales simultaneously will bring enormous storage and computing pressure to spacecraft equipment.
@@
In contrast, arbitrary-scale SR uses a single model to process LR images at arbitrary scales, which is more suitable for the practical image SR scenarios. Most current arbitrary-scale SR methods, such as LIIF [14] and LTE [15], are based on implicit neural representation (INR), and use implicit neural function to replace the upsampling module in traditional image SR methods. However, these methods are aimed at general images and failed to fully consider the characteristics of spacecraft images.
@@
As shown in Fig. 1, spacecraft images acquired from the outer space environment, and mostly with a large black space background. The existing arbitrary-scale SR methods regard each region in spacecraft images equally, overlooking the differences in features between the spacecraft core region and the background region [16]. As a result, these models are prone to falling into the local suboptimal solutions or even introducing noise to the black background region. In addition, the SR model brings more significant improvements to the spacecraft core region than to the background region. This limitation makes it difficult for existing arbitrary-scale SR methods to be effectively applied to spacecraft image SR tasks. Therefore, how to better focus on restoring the spacecraft core region is the key to improving the performance of the spacecraft image SR task.
In this paper, we propose a salient region-guided spacecraft image arbitrary-scale super-resolution network (SGSASR). Specifically, we design an arbitrary-scale image SR framework based on latent modulated function (LMF), which generates latent modulation guided by the spacecraft core salient region, and then decodes the latent modulations to generate spacecraft SR images of different scales according to requirements. Furthermore, we design a spacecraft core region recognition block (SCRRB) to identify the core salient region in spacecraft LR images, allowing the network to focus on restoring the region. Finally, we propose an adaptive-weighted feature fusion enhancement mechanism (AFFEM), which selectively aggregates the spacecraft core region features with general image features to enhance the response of the core salient region.
@@
  • Extensive experiments demonstrate that our SGSASR not only achieves state-of-the-art performance on spacecraft image arbitrary-scale SR, but also achieves competitive performance compared to the fixed-scale SR method.
This paper is organized as follows. Section II introduces the related works. Section III describes our proposed method in detail. Section IV presents the experimental results and analysis. Section V concludes the paper.
II Related Work
II-A Spacecraft Image Super-Resolution
Spacecraft image super-resolution (SR) task aims to synthesize high-resolution (HR) spacecraft images from low-resolution (LR) spacecraft images. The early spacecraft image SR methods typically required manually designed prior features [17, 18, 19]. In recent years, deep learning-based methods have gradually been introduced into the spacecraft image SR task. Zhang et al. [20] propose a spacecraft image SR method based on the deep recursive convolutional network and fine-tune it on the simulated space object dataset. Yang et al. [11] propose a unified GAN method for spacecraft image SR and deblurring, which restores texture information by designing symmetric downsampling and upsampling modules. Shi et al. [21] introduce the dual regression and deformable convolutional attention mechanism to better extract the high-frequency features of spacecraft LR images. Yang et al. [12] propose a spacecraft ISAR image SR method based on non-local sparse attention, which can capture long-range self-similarity in spacecraft images. Li et al. [22] present a non-local scene awareness network (NLSAN), which incorporates a non-local interpolation approach to capture global textures in spacecraft images. Gao et al. [23] propose a sparse attention and selective cross-fusion network (SSNet) for spacecraft ISAR image SR, which introduces the top-k selection operator to enhance the distinction between background and spacecraft information.
II-B Single Image Super-Resolution
Single Image Super-Resolution (SISR) is a classic image SR task, and a large number of deep learning-based SISR methods have been proposed in recent years [24, 25, 26]. These methods can be classified into two main categories: CNN-based and Transformer-based.
@@ II-B Single Image Super-Resolution
Transformer is also introduced into SISR in recent years. IPT [34] is the first Transformer-based SISR method, which adapts to different image processing tasks by pre-training on the ImageNet benchmark. Zhang et al. [35] introduce an efficient long-range attention network (ELAN) that combines shift convolution and a group-wise multi-scale self-attention module to reduce computational complexity while exploiting long-range image dependencies. Zhao et al. [36] propose a spherical space feature decomposition network that effectively separates and aligns modality-specific features by contrastive learning in the spherical space. Li et al. [37] propose a lightweight SISR framework (SRConvNet) that integrates the global modeling strengths of transformers with the efficiency of CNNs by introducing Fourier-modulated attention and dynamic mixing layer. Xiao et al. [38] creatively introduce the top-k token selective mechanism, dynamically selecting the top-k keys in terms of score ranking for each query to obtain the most crucial token.
II-C Arbitrary-Scale Super-Resolution
@@
Arbitrary-scale super-resolution is a task that uses a single model to achieve SR at any scale, so it is more practical. Hu et al. [39] propose MetaSR, which uses meta-module to dynamically predict the weights of large-scale filters based on input scale. Chen et al. [14] propose the local implicit image function (LIIF) inspired by implicit neural representation (INR), which predicts RGB values at arbitrary coordinates using local deep features. Subsequently, INR becomes the main paradigm for arbitrary-scale SR tasks [40, 41, 42]. Lee et al. [15] enhance implicit neural functions by capturing fine image textures in the 2D Fourier space. He et al. [43] introduce a computationally optimal paradigm for continuous image representation in arbitrary-scale SR by decoupling HR decoding into a shared latent space and a lightweight rendering space. Zhao et al. [44] propose SLFAM and LFAUM that enable fixed-scale SR networks to effectively handle arbitrary scales by incorporating both scale factors and local feature diversity. Zhu et al. [45] propose the multi-scale implicit transformer (MSIT), which introduces multi-scale characteristics into arbitrary-scale SR for the first time, enhancing latent code adaptability across magnification factors.
III Methodology
@@ III-A Overall Architecture
Figure 3: (a) The architecture of spacecraft image feature extraction enhancement module (SFEEM). (b) Spacecraft core region recognition block (SCRRB). (c) Nonlinear activation free block (NAFBlock). (d) Simplified channel attention (SCA). (e) Simple gate. (f) Adaptive-weighted feature fusion enhancement mechanism (AFFEM).
III-B Spacecraft Image Feature Extraction Enhancement Module (SFEEM)
@@
III-B1 Spacecraft Core Region Recognition Block (SCRRB)
In order to identify the spacecraft core region in the image, we need to perform saliency object detection on the input spacecraft image. However, directly inserting a saliency detection module into the network structure cannot guarantee detection accuracy and increases the additional computational burden. Therefore, we propose to introduce a pre-trained object detection model and construct SCRRB based on it. The architecture of SCRRB is shown in Fig. 3 (b).
@@ III-B1 Spacecraft Core Region Recognition Block (SCRRB)
III-B2 Nonlinear Activation Free Block (NAFBlock)
In order to improve computational efficiency, we utilize NAFBlock [47] as the base block of the encoder-decoder in SFEEM. NAFBlock introduces the simple gate function to replace the nonlinear activation function, as shown in Fig. 3 (c). Formally, the detailed process of NAFBlock is represented as follows:
@@ III-B2 Nonlinear Activation Free Block (NAFBlock)
III-B3 Adaptive-Weighted Feature Fusion Enhancement Mechanism (AFFEM)
The traditional feature fusion methods such as summation and concatenation lack flexible expression, which limits the performance of feature fusion. Therefore, we propose AFFEM as shown in Fig. 3 (f), which achieves adaptive feature fusion enhancement by dynamic weight parameters. Formally, the process of our AFFEM can be defined as:
@@ III-B3 Adaptive-Weighted Feature Fusion Enhancement Mechanism (AFFEM)
III-C Latent Modulation-Based arbitrary-scale super-resolution Module (LMASSRM)
@@
III-C1 Latent Modulation Generation Block (LMGB)
In LMGB, we use latent MLP to generate latent modulation $m$ according to the input latent code $z$, and compress the high-dimensional latent code $z$ into the low-dimensional latent code $z_{c}$. The latent modulation encodes the input features and provides dynamic parameters for subsequent restoration. The latent modulation $m$ consists of the FiLM layers [48], each of which includes the scale modulation and the shift modulation.
@@ III-C1 Latent Modulation Generation Block (LMGB)
TABLE I: Quantitative results on spacecraft ISAR image dataset. The best and the second best results are marked in bold and underline.
Methods $\times$4 $\times$6 $\times$8 $\times$12
[…]
SSRN [12] 32.93 0.9847 - - 28.64 0.9730 - -
SSNet [23] 32.23 0.9828 - - 27.91 0.9627 - -
VapSR [31] 32.32 0.9825 - - 28.30 0.9667 - -
LKFN [33] 31.77 0.9817 - - 27.94 0.9640 - -
ELAN [35] 31.06 0.9678 - - 28.06 0.9340 - -
SRConvNet [37] 32.37 0.9822 - - 28.17 0.9634 - -
LIIF [14] 24.88 0.9588 25.96 0.9563 25.03 0.9583 23.74 0.9501
LTE [15] 31.01 0.9737 29.17 0.9662 28.17 0.9638 27.12 0.9550
LM-LTE [43] 31.73 0.9758 29.54 0.9685 28.51 0.9657 27.27 0.9570
SGSASR(Our approach) 34.07 0.9823 30.24 0.9758 29.03 0.9723 28.22 0.9700
III-C2 Arbitrary Scale Restoration Block (ASRB)
@@
Figure 5: Visual comparison for fixed-scale SR methods on spacecraft ISAR image dataset. (a): ×$\times 4$. (b): ×$\times 8$.
Figure 6: Visual comparison for arbitrary-scale SR methods on spacecraft ISAR image dataset. (a): ×$\times 4$. (b): ×$\times 6$. (c): ×$\times 8$. (d): ×$\times 12$.
IV Experiment
@@
IV-A Experimental Settings
IV-A1 Datasets
In this paper, we conduct experiments on two spacecraft image datasets. The first is a real spacecraft ISAR image dataset, and the second is the publicly available spacecraft optical image dataset SPEED [49].
The real spacecraft ISAR image dataset consists of real ISAR data collected during the in-orbit flight of a certain type of non-cooperative spacecraft. It contains a total of 5,250 images, with 4,800 used for training and 450 for evaluation.
@@ IV-A1 Datasets
IV-A2 Evaluation Metrics
We use two popular image quality evaluation metrics: peak signal-to-noise ratio (PSNR) and structural similarity (SSIM). The higher values of these two metrics indicate the better performance of the method.
IV-A3 Implementation Details
@@
IV-B Comparison Results on Spacecraft ISAR Image Dataset
In this section, we compare SGSASR with six fixed-scale SR methods (SSRN [12] and SSNet [23], which are spacecraft image SR methods, as well as VapSR [31], LKFN [33], ELAN [35], and SRConvNet [37]), and three arbitrary-scale SR methods (LIIF [14], LTE [15] , and LM-LTE [43]) on the spacecraft ISAR image dataset. All comparison methods are retrained to ensure fairness and effectiveness. Table I presents the quantitative comparison results. We can observe that SGSASR significantly outperforms the fixed-scale SR methods in terms of PSNR performance on spacecraft ISAR image dataset. Specifically, compared with SSRN, SGSASR respectively achieves improvements of 1.14 dB and 0.39 dB at two SR scales ($\times$4, $\times$8). We can also observe that SGSASR consistently achieves superior PSNR performance at various SR scales compared with other arbitrary-scale SR methods on spacecraft ISAR image dataset. Specifically, compared with LM-LTE, SGSASR respectively achieves improvements of 2.36 dB, 0.70 dB, 0.52 dB and 0.95 dB at four SR scales ($\times$4, $\times$6, $\times$8, $\times$12).

### v2

# Towards Arbitrary-Scale Spacecraft Image Super-Resolution via Salient Region-Guidance

@@ Abstract
Spacecraft image super-resolution seeks to enhance low-resolution spacecraft images into high-resolution ones. Although existing arbitrary-scale super-resolution methods perform well on general images, they tend to overlook the difference in features between the spacecraft core region and the large black space background, introducing irrelevant noise. In this paper, we propose an efficient salient region-guided spacecraft image arbitrary-scale super-resolution network (SGSASR), which uses features from the spacecraft core salient regions to guide latent modulation and achieve arbitrary-scale super-resolution. Specifically, we design a spacecraft core region recognition block (SCRRB) that identifies the core salient regions in spacecraft images using a pre-trained saliency detection model. Furthermore, we present an adaptive-weighted feature fusion enhancement mechanism (AFFEM) to selectively aggregate the spacecraft core region features with general image features by dynamic weight parameter to enhance the response of the core salient regions. Experimental results on spacecraft radar image dataset and optical image dataset demonstrate that the proposed SGSASR outperforms state-of-the-art approaches. The codes are available at: https://github.com/shenduke/SGSASR.
keywords:
@@
Arbitrary-scale super-resolution, Latent modulation, Salient region-guided
journal: Pattern Recognition
[…]
[a]organization=School of Artificial Intelligence, Beijing Normal University,addressline=No.19, Xinjiekouwai St., city=Beijing, postcode=100875, country=China
1 Introduction
@@
Spacecraft image is the on-orbit spacecraft target image obtained through radar or optical technology. The clear and high-quality spacecraft image is the fundamental assurance for successfully completing many space missions such as spacecraft pose estimation [1, 2], spacecraft control [3, 4] and spacecraft 3D reconstruction [5, 6]. However, due to the large imaging distance, the harsh space lighting environment and sensor interference, it is difficult to obtain high-resolution (HR) spacecraft image [7].
@@
Spacecraft image super-resolution (SR) aims to restore HR spacecraft images from low-resolution (LR) spacecraft images. Although many spacecraft image SR methods have been proposed [8, 9], they typically learn upsampling functions for a specific scale factor and require training different models for different SR scales [10]. When there is a lack of suitable scale training data or the degradation scale of images is unfixed, these methods yield unsatisfactory performance. In practical aerospace application scenarios, it is relatively difficult to obtain real spacecraft HR images to construct training datasets. The degradation scale of different images in the same spacecraft image dataset is also unfixed because of the complex space environmental conditions. In addition, storing a large number of models with different super-resolution scales simultaneously will bring enormous storage and computational pressure to spacecraft equipment.
@@
In contrast, arbitrary-scale SR uses a single model to process LR images at arbitrary scales, which is more suitable for practical image SR scenarios. Most current arbitrary-scale SR methods, such as LIIF [11] and LTE [12], are based on implicit neural representation (INR), and use the implicit neural function to replace the upsampling module in traditional image SR methods. However, these methods are aimed at general images and failed to fully consider the characteristics of spacecraft images.
@@
As shown in Fig. 1, spacecraft images are acquired from the outer space environment, and mostly with a large black space background. The existing arbitrary-scale SR methods regard each region in spacecraft images equally, overlooking the differences in features between the spacecraft core region and the background region [13]. As a result, these models are prone to falling into local suboptimal solutions or even introducing noise to the black background region. This limitation makes it difficult for existing arbitrary-scale SR methods to be effectively applied to spacecraft image SR tasks. Therefore, how to better focus on restoring the spacecraft core region is the key to improving the performance of the spacecraft image SR task.
In addition, spacecraft image application scenarios, such as spacecraft control and autonomous navigation, require high real-time performance to ensure timely decision-making and reliable execution of critical operations.
To address these challenges, we propose an efficient salient region-guided spacecraft image arbitrary-scale super-resolution network (SGSASR) in this paper. Specifically, we design an arbitrary-scale image SR framework based on latent modulated function (LMF), which generates latent modulation guided by the spacecraft core salient region, and then decodes the latent modulations to generate spacecraft SR images of different scales according to requirements. Furthermore, we design a spacecraft core region recognition block (SCRRB) to identify the core salient region in spacecraft LR images, allowing the network to focus on restoring the region. Finally, we propose an adaptive-weighted feature fusion enhancement mechanism (AFFEM), which selectively aggregates the spacecraft core region features with general image features to enhance the response of the core salient region.
@@
  • 4.
    Extensive experiments demonstrate that our SGSASR not only achieves state-of-the-art performance on spacecraft image arbitrary-scale SR while maintaining high computational efficiency, but also achieves competitive performance compared to fixed-scale SR methods.
This paper is organized as follows. Section 2 introduces the related works. Section 3 describes our proposed method in detail. Section 4 presents the experimental results and analysis. Section 5 concludes the paper.
2 Related Work
2.1 Single Image Super-Resolution
Single Image Super-Resolution (SISR) is a classic image SR task, and a large number of deep learning-based SISR methods have been proposed in recent years [14, 15]. These methods can be classified into two main categories: CNN-based and Transformer-based.
@@ 2.1 Single Image Super-Resolution
Transformer has also been introduced into SISR in recent years. IPT [23] is the first Transformer-based SISR method, which adapts to different image processing tasks by pre-training on the ImageNet benchmark. Liang et al. [24] propose SwinIR, an image SR method based on the Swin Transformer, which combines shallow feature extraction, deep feature extraction with residual Swin Transformer blocks, and high-quality image reconstruction. Zhang et al. [25] introduce an efficient long-range attention network (ELAN) that combines shift convolution and a group-wise multi-scale self-attention module to reduce computational complexity while exploiting long-range image dependencies. Zhao et al. [26] propose a spherical space feature decomposition network that effectively separates and aligns modality-specific features by contrastive learning in the spherical space. Xie et al. [27] propose a Multi-Range Attention Transformer (MAT), which flexibly integrates attention across diverse spatial extents using dilation-based operations. Li et al. [28] propose a lightweight SISR framework (SRConvNet) that integrates the global modeling strengths of transformers with the efficiency of CNNs by introducing Fourier-modulated attention and dynamic mixing layer. Xiao et al. [29] creatively introduce the top-k token selective mechanism, dynamically selecting the top-k keys in terms of score ranking for each query to obtain the most crucial token.
2.2 Spacecraft Image Super-Resolution
The spacecraft image super-resolution (SR) task aims to generate high-resolution (HR) spacecraft images from low-resolution (LR) inputs. The early spacecraft image SR methods typically required manually designed prior features [30, 31]. In recent years, deep learning-based methods have gradually been introduced into the spacecraft image SR task. Zhang et al. [32] propose a spacecraft image SR method based on the deep recursive convolutional network and fine-tune it on the simulated space object dataset. Yang et al. [8] propose a unified GAN method for spacecraft image SR and deblurring, which restores texture information by designing symmetric downsampling and upsampling modules. Shi et al. [33] introduce the dual regression and deformable convolutional attention mechanism to better extract the high-frequency features of spacecraft LR images. Yang et al. [9] propose a spacecraft ISAR image SR method based on non-local sparse attention, which can capture long-range self-similarity in spacecraft images. Li et al. [34] present a non-local scene awareness network (NLSAN), which incorporates a non-local interpolation approach to capture global textures in spacecraft images. Gao et al. [35] propose a sparse attention and selective cross-fusion network (SSNet) for spacecraft ISAR image SR, which introduces the top-k selection operator to enhance the distinction between background and spacecraft information.
2.3 Arbitrary-Scale Super-Resolution
@@
Arbitrary-scale super-resolution is a task that uses a single model to achieve SR at any scale, so it is more practical. Hu et al. [36] propose MetaSR, which uses meta-module to dynamically predict the weights of large-scale filters based on input scale. Chen et al. [11] propose the local implicit image function (LIIF) inspired by implicit neural representation (INR), which predicts RGB values at arbitrary coordinates using local deep features. Subsequently, INR becomes the main paradigm for arbitrary-scale SR tasks [37, 38]. Lee et al. [12] enhance implicit neural functions by capturing fine image textures in the 2D Fourier space. He et al. [39] introduce a computationally optimal paradigm for continuous image representation in arbitrary-scale SR by decoupling HR decoding into a shared latent space and a lightweight rendering space. Zhao et al. [40] propose SLFAM and LFAUM that enable fixed-scale SR networks to effectively handle arbitrary scales by incorporating both scale factors and local feature diversity. Wei et al. [41] propose a Dynamic Implicit Network (DINet) for multi-contrast MRI arbitrary-scale SR, which uses the scale-adaptive dynamic convolution to learn multi-contrast feature representations. Zhu et al. [42] propose the multi-scale implicit transformer (MSIT), which introduces multi-scale characteristics into arbitrary-scale SR for the first time, enhancing latent code adaptability across magnification factors. Zhang et al. [43] propose DASU-Net and its fast version FDASU-Net, which is a joint depth map super-resolution and high–low frequency decomposition framework capable of arbitrary-scale up-sampling and robust cross-modal feature fusion.
3 Methodology
@@ 3.1 Overall Architecture
Figure 3: The architecture of spacecraft image feature extraction enhancement module (SFEEM).
3.2 Spacecraft Image Feature Extraction Enhancement Module (SFEEM)
@@
Figure 4: (a) Spacecraft core region recognition block (SCRRB). (b) Nonlinear activation free block (NAFBlock). (c) Simplified channel attention (SCA). (d) Simple gate. (e) Adaptive-weighted feature fusion enhancement mechanism (AFFEM).
3.2.1 Spacecraft Core Region Recognition Block (SCRRB)
In order to identify the spacecraft core region in the image, we need to perform saliency object detection on the input spacecraft image. However, directly inserting a saliency detection module into the network structure cannot guarantee detection accuracy and increases the additional computational burden. Therefore, we propose to introduce a pre-trained object detection model and construct SCRRB based on it. The architecture of SCRRB is shown in Fig. 4 (a).
@@ 3.2.1 Spacecraft Core Region Recognition Block (SCRRB)
3.2.2 Nonlinear Activation Free Block (NAFBlock)
In order to improve computational efficiency, we utilize NAFBlock [45] as the base block of the encoder-decoder in SFEEM. NAFBlock introduces the simple gate function to replace the nonlinear activation function, as shown in Fig. 4 (b). Formally, the detailed process of NAFBlock is represented as follows:
@@ 3.2.2 Nonlinear Activation Free Block (NAFBlock)
3.2.3 Adaptive-Weighted Feature Fusion Enhancement Mechanism (AFFEM)
The traditional feature fusion methods such as summation and concatenation lack flexible expression, which limits the performance of feature fusion. Therefore, we propose AFFEM as shown in Fig. 4 (e), which achieves adaptive feature fusion enhancement by dynamic weight parameters. Formally, the process of our AFFEM can be defined as:
@@ 3.2.3 Adaptive-Weighted Feature Fusion Enhancement Mechanism (AFFEM)
3.3 Latent Modulation-Based arbitrary-scale super-resolution Module (LMASSRM)
@@
3.3.1 Latent Modulation Generation Block (LMGB)
In LMGB, we use latent MLP to generate latent modulation $m$ according to the input latent code $z$, and compress the high-dimensional latent code $z$ into the low-dimensional latent code $z_{c}$. The linear layers in the latent MLP are modulated using the FiLM mechanism [46], which includes the scale modulation and the shift modulation. The latent modulation $m$ encodes the input features and provides dynamic parameters for subsequent restoration.
@@ 3.3.1 Latent Modulation Generation Block (LMGB)
3.3.2 Arbitrary Scale Restoration Block (ASRB)
@@
4 Experiment
@@
4.1 Experimental Settings
4.1.1 Datasets
In this paper, we conduct experiments on two spacecraft image datasets. The first is a real spacecraft radar image dataset, and the second is the publicly available spacecraft optical image dataset SPEED [47].
The real spacecraft radar image dataset consists of real inverse synthetic aperture radar (ISAR) images collected during the in-orbit flight of a certain type of non-cooperative spacecraft. It contains a total of 5,250 images, with 4,800 used for training and 450 for evaluation.
@@ 4.1.1 Datasets
4.1.2 Evaluation Metrics
We use two popular image quality evaluation metrics: peak signal-to-noise ratio (PSNR) and structural similarity (SSIM). The higher values of these two metrics indicate the better performance of the method. Furthermore, we use Floating Point Operations (FLOPs) to evaluate the computational complexity, the lower value of the FLOPs indicates the lower the computational complexity and the higher the computing efficiency of the method.
4.1.3 Implementation Details
@@
Table 1: Quantitative comparisons with other fixed-scale SR methods on spacecraft radar image dataset. The best and the second best results are marked in bold and underline.
Methods $\times$4 $\times$8
$uparrow PSNR\uparrow SSIM\uparrow$
SSRN [9] 32.93 0.9847 28.64 0.9730
SSNet [35] 32.23 0.9828 27.91 0.9627
VapSR [20] 32.32 0.9825 28.30 0.9667
LKFN [22] 31.77 0.9817 27.94 0.9640
ELAN [25] 31.06 0.9678 28.06 0.9340
MAT [27] 32.07 0.9781 28.56 0.9568
SRConvNet [28] 32.37 0.9822 28.17 0.9634
SGSASR(Ours) 34.07 0.9823 29.03 0.9723
Table 2: Quantitative comparisons with other arbitrary-scale SR methods on spacecraft radar image dataset. The best and the second best results are marked in bold and underline.
[…]
$uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$
LIIF [11] 33.51 0.9826 30.15 0.9765 28.95 0.9730 28.05 0.9707 160.64G
LM-LIIF [39] 34.17 0.9825 30.19 0.9757 28.99 0.9721 27.99 0.9699 20.99G
LM-LTE [39] 33.03 0.9819 30.07 0.9755 28.88 0.9718 28.04 0.9685 21.55G
SGSASR(Ours) 34.07 0.9823 30.24 0.9758 29.03 0.9723 28.22 0.9700 4.37G
Figure 6: Visual comparison for fixed-scale SR methods on spacecraft radar image dataset. (a): $\times 4$. (b): $\times 8$.
Figure 7: Visual comparison for arbitrary-scale SR methods on spacecraft radar image dataset. (a): $\times 4$. (b): $\times 6$. (c): $\times 8$. (d): $\times 12$.
4.2 Comparison Results on Spacecraft Radar Image Dataset
In this section, we compare SGSASR with seven fixed-scale SR methods (SSRN [9] and SSNet [35], which are spacecraft image SR methods, as well as VapSR [20], LKFN [22], ELAN [25], MAT [27] and SRConvNet [28]), and three arbitrary-scale SR methods (LIIF [11], LM-LIIF [39], and LM-LTE [39]) on the spacecraft radar image dataset. All comparison methods are retrained to ensure fairness and effectiveness. Table 1 presents the quantitative comparison results with fixed-scale SR methods. Table 2 presents the quantitative comparison results with arbitrary-scale SR methods. We can observe that SGSASR significantly outperforms the fixed-scale SR methods in terms of PSNR performance on the spacecraft radar image dataset. Specifically, compared with SSRN, SGSASR achieves improvements of 1.14 dB and 0.39 dB at two SR scales ($\times$4, $\times$8), respectively. We can also observe that SGSASR consistently achieves superior PSNR performance at various SR scales on the spacecraft radar image dataset while maintaining lower computational costs compared with other arbitrary-scale SR methods.